OpenStudy (anonymous):

Write each of the following square roots in the form a square root b where a and b are integers and b has the least value possible: a. sq. rt 360 b. sq. rt 40 c. sq. rt 240

OpenStudy (anonymous):

\[A. \sqrt{360} = \sqrt[?]{?}, B. \sqrt{40} = \sqrt[?]{?}, C. \sqrt{240} = \sqrt[?]{?}\]

OpenStudy (anonymous):

360=36x10 so sqrt(360)=sqrt(36x10)=6sqrt(10) do the rest..

OpenStudy (anonymous):

Okay kool I guess I was on the right track thank you

OpenStudy (anonymous):

so isthe sq. rt of 240 2 sq. rt. 15

OpenStudy (asnaseer):

\[\sqrt{240}=\sqrt{16*15}=4\sqrt{15}\]